OpenStudy (anonymous):

Two sides of a triangle have lengths 10 and 18. Which inequalities describe the values that possible lengths for the third side? I'll tell you all the possible answers but I need to know how to solve which one

OpenStudy (confluxepic):

The third side has to be less than the sum of the other two sides.

OpenStudy (anonymous):

if a,b and c are the sides of a triangle, then |a-b|<c<a+b

OpenStudy (confluxepic):

c<a+b

OpenStudy (anonymous):

\[A) 8 \le x \le 28 B) 8 < x < 28 C) 10 < x < 18 D) 10 \le x \le 18\]

OpenStudy (anonymous):

all possible answers

OpenStudy (anonymous):

B

OpenStudy (anonymous):

So I got 18 math mind

OpenStudy (anonymous):

say a = 10 and b=28

OpenStudy (confluxepic):

x has to be less than 10+18 which is 28. It can not be equal to. So B is the only answer.

OpenStudy (anonymous):

|a-b|<c<a+b |10-18|<c<10+18 |-8|<c<28 8<c<28

OpenStudy (anonymous):

I see now thank you both!

OpenStudy (anonymous):

no problem ;)

OpenStudy (anonymous):

Yeah! i meant 18